OpenStudy (anonymous):

The equation 3x2 – 6y2 = 6 represents which conic section: -circle -parabola -hyperbola -ellipse

OpenStudy (solomonzelman):

\[ 3x^2 – 6y^2 = 6\] \[ \frac{x^2}{2} – \frac{y^2}{1} = 1\]

OpenStudy (solomonzelman):

So, what do you think ?

OpenStudy (anonymous):

a hyperbola